You are given the following information of a company that produces cola fizzy drinks. The

company states that the mean caffeine content per bottle of a fizzy drink is 40 mg with a

standard deviation of 7.5 mg. The quality controller is convinced that it is lower. A sample of

30 randomly drawn bottles has a mean caffeine content of 39.2 mg. Can the quality

controller reject the claim? Conduct a hypothesis test at 0.05 level of significance.


1
Expert's answer
2022-06-07T09:26:21-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge40"

"H_1:\mu<40"

This corresponds to a left-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a left-tailed test is "z_c = -1.6449."

The rejection region for this left-tailed test is "R = \{z:z<-1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{39.2-40}{7.5/\sqrt{30}}=-0.5842"

Since it is observed that "z=-0.5842>-1.6449=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=P(z<-0.5842)= 0.279543," and since "p= 0.279543>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is less than 40, at the "\alpha = 0.05" significance level.